# Spectroscopy of $N=50$ isotones with the valence-space density matrix renormalization group

### v1

@@ 5 Low-lying spectroscopy of $N=50$ isotones
With the performance of the VS-DMRG being validated for spectroscopic observables, we now turn to studying the low-lying spectroscopy of $N=50$ isotones from calcium ($Z=20$) to zink ($Z=30$). Due to their neutron-rich character, except for ${}^{78}$i and ${}^{80}$n, no spectroscopy data is available Taniuchi et al. (2019). Hence we compare our results to state-of-the-art shell-model calculations by Nowacki et al. Nowacki et al. (2016). They observed a rapid transition from single-particle-like excitations in ${}^{78}$i to collective rotational excitations in ${}^{74}$r. Rotational structures can be approximately extracted by comparing the level spacings of the low-lying spectrum to that of a rigid rotor, $E^{\star}(J)\sim J(J+1)$ Bohr and Mottelson (1998). For the case of the lowest $4^{+}$ and $2^{+}$ state the ratio of the energies $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}$ is given in the top panel of Fig. 3. For a perfect rotor the expected value is $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}{}=3.33$, whereas for single-particle-like excitations the ratio will be lower. Starting from doubly magic ${}^{78}$i there is a rapid transition to the rotational spectrum of ${}^{74}$r with a ratio of $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}{}=2.83$. Our VS-DMRG predictions qualitatively match the shell-model results from Ref. Nowacki et al. (2016) with noticeable deviation in ${}^{72}$i. Still there is a clear transition towards the rigid-rotor value in ${}^{74}$r, which suggests that our ab initio calculations qualitatively capture this deformation physics.
@@
Figure 3: Top panel: Ratio of $4^{+}$ to $2^{+}$ excitation energies of $N=50$ isotones for our VS-DMRG results and for the phenomenological shell-model calculations from Nowacki et al. Nowacki et al. (2016), in comparison to a rotational and harmonic spectrum. Middle panel: Excitation energies and $B(E2)$ transition strengths for the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states from our VS-DMRG calculations at bond dimension $M=10240$. Bottom panel: $T_{\text{max}}=0$ component of the ground-state wave function (orange fraction).
@@
Moreover, we show in Fig. 3 the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states with the allowed $B(E2)$ transitions. The $2^{+}$ state in ${}^{78}$i at $E^{\star}\approx 3.0\,\text{MeV}$ is significantly higher in energy due to the proton shell closure at $Z=28$ and we observe a rapid decrease towards a mid-shell minimum at chromium ($Z=24$) with $E^{\star}\approx 0.5\,\text{MeV}$ and another increase towards calcium. The calculated $B(E2)$ values show a maximum for ${}^{74}$r for the $2^{+}\rightarrow 0^{+}$ transition that is characteristic for collective rotational excitations and a signature of nuclear deformation. In Ref. Nowacki et al. (2016) this was explicitly confirmed by performing a symmetry-broken and -restored mean-field calculations revealing a pronounced prolate minimum for ${}^{74}$r with axial deformation parameter $\beta\approx 0.35$.
@@
We further show the fraction of the $T_{\text{max}}=0$ component of the ground-state wave function in the bottom panel of Fig. 3. While in the case of ${}^{78}$i this component is given by $|c_{0}|^{2}\approx 0.48$ indicating a dominant occupation of the energetically lowest configurations in the ground state, this changes drastically towards mid-shell ${}^{74}$r with very small $|c_{0}|^{2}\approx 0.03$. Similar observation hold to a lesser extent for the neighbouring nuclei ${}^{76}$e ($|c_{0}|^{2}\approx 0.28)$ and ${}^{72}$i ($|c_{0}|^{2}\approx 0.19$).
@@ 6 Quantum information analysis of ground states
Figure 4: Single-orbital entropies for the ground state of ${}^{74}$r in the VS-DMRG, in comparison with increasing CI particle-hole truncations. The corresponding total entropies are given in the legend.
We finally turn to a discussion of orbital entanglement for the strongly correlated ${}^{74}$r nucleus using established QI measures. Many-body correlation effects can be explicitly linked to orbital entanglement between single-particle states for a given nuclear wave function Legeza et al. (2015). We therefore, define the subsystem’s reduced density matrix $A\subset\mathcal{H}_{1}$ via $\rho_{A}=\text{Tr}_{B}\,\rho_{AB}$, where $A\cup B=\mathcal{H}_{1}$. Setting $A=\{p\}$ we introduce the single-orbital entropy as $s_{i}=-\text{Tr}\,\rho\,\log\,\rho$ that is directly linked to the occupation numbers $n_{p}$ via $s_{p}=-n_{p}\,\log\,n_{p}-\bar{n}_{p}\,\log\,\bar{n}_{p}$ where $\bar{n}_{p}\equiv 1-n_{p}$. Therefore, single-orbital entropies can be straightforwardly evaluated in the DMRG and CI frameworks. Our previous work showed that the orbital entropies can be linked to the closed-shell character of a nucleus Tichai et al. (2023).
@@ 7 Conclusion and outlook
Finally our methodology was applied to $N=50$ isotones to obtain an ab initio description of transitional nuclei at the $N=50$ shell closure. While our predictions qualitatively agree with previous state-of-the-art shell-model calculations, we fail to reproduce the deformed excited-state band in ${}^{78}$i, hinting at increased many-body uncertainties or interaction deficiencies.
While the computational demands of the VS-DMRG implementation are tractable in this work, we envision more advanced schemes to cope with computational challenges emerging in larger valence spaces. This can be either achieved by a symmetry-adapted formulation of the DMRG explicitly accounting for non-Abelian SU(2) symmetries (see Ref. McCulloch and Gulácsi (2002)), or by a more careful subspace selection using restricted-active-space formulations, where the coupling to weakly entangled high-energy orbits is approximated by limiting oneself to low-rank CI excitations Barcza et al. (2022); Friesecke et al. (2023). This work paves the way for accessing configuration spaces well above current shell-model capacities. These arise for heavy nuclei and nuclei far away from shell closures, and will be particularly important for deriving ab initio constraints on heavy neutron-rich nuclei.
@@ Acknowledgements
The work of A.T., T.M., and A.S. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842). M.A.W. was supported by the Bolyai Research Scholarship of the Hungarian Academy of Sciences. Ö.L. and G.Z. were supported by the Quantum Information National Laboratory of Hungary and by the Hungarian National Research Development and Innovation Office (NKFIH) through Grants Nos. K134983, SNN139581, and TKP2021-NVA-04. Ö.L. recieved further support from the Hans Fischer Senior Fellowship programme funded by the Technical University of Munich – Institute for Advanced Study and from the Center for Scalable and Predictive methods for Excitation and Correlated phenomena (SPEC), funded as part of the Computational Chemical Sciences Program by the U.S. Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences, and Biosciences at Pacific Northwest National Laboratory. Computations were in part performed with an allocation of computing resources at the Jülich Supercomputing Center. We have used the NuHamil Miyagi (2023), imsrg++ Stroberg (2019), and KSHELL Shimizu et al. (2019) codes for the 1.8/2.0 (EM) NN+3N matrix elements generation, VS-IMSRG decoupling, and CI diagonalization, respectively.

### v2

@@ 5 Low-lying spectroscopy of $N=50$ isotones
With the performance of the VS-DMRG being validated for spectroscopic observables, we now turn to studying the low-lying spectroscopy of $N=50$ isotones from calcium ($Z=20$) to zinc ($Z=30$). Due to their neutron-rich character, except for 78Ni and 80Zn, no spectroscopy data is available [44]. Hence we compare our results to state-of-the-art shell-model calculations by Nowacki et al. [45]. They observed a rapid transition from single-particle-like excitations in 78Ni to collective rotational excitations in 74Cr. Rotational structures can be approximately extracted by comparing the level spacings of the low-lying spectrum to that of a rigid rotor, $E^{\star}(J)\sim J(J+1)$ [68]. For the case of the lowest $4^{+}$ and $2^{+}$ state the ratio of the energies $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}$ is given in the top panel of Fig. 3. For a perfect rotor the expected value is $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}{}=3.33$, whereas for single-particle-like excitations the ratio will be lower. Starting from doubly magic 78Ni there is a rapid transition to the rotational spectrum of 74Cr with a ratio of $E^{\star}_{4^{+}}/E^{\star}_{2^{+}}{}=2.83$. Our VS-DMRG predictions qualitatively match the shell-model results from Ref. [45] with noticeable deviation in 72Ti. Still there is a clear transition towards the rigid-rotor value in 74Cr, which suggests that our ab initio calculations qualitatively capture this deformation physics.
@@
Figure 3: Top panel: Ratio of $4^{+}$ to $2^{+}$ excitation energies of $N=50$ isotones for our VS-DMRG results and for the phenomenological shell-model calculations from Nowacki et al. [45], in comparison to a rotational and single-particle-like (“harmonic”) spectrum. Middle panel: Excitation energies and $B(E2)$ transition strengths in units of e2 fm4 for the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states from our VS-DMRG calculations at bond dimension $M=10240$. Bottom panel: $T_{\text{max}}=0$ component of the ground-state wave function (orange fraction).
@@
Moreover, we show in Fig. 3 the lowest $0^{+}$, $2^{+}$, and $4^{+}$ states with the allowed $B(E2)$ transitions. The $2^{+}$ state in 78Ni at $E^{\star}\approx 3.0\,\text{MeV}$ is significantly higher in energy due to the proton shell closure at $Z=28$ and we observe a rapid decrease towards a mid-shell minimum at chromium ($Z=24$) with $E^{\star}\approx 0.5\,\text{MeV}$ and another increase towards calcium. The calculated $B(E2)$ values show a maximum for 74Cr for the $2^{+}\rightarrow 0^{+}$ transition that is characteristic for collective rotational excitations and a signature of nuclear deformation. In Ref. [45] this was explicitly confirmed by performing a symmetry-broken and -restored mean-field calculations revealing a pronounced prolate minimum for 74Cr with axial deformation parameter $\beta\approx 0.35$.
@@
We further show the fraction of the $T_{\text{max}}=0$ component of the ground-state wave function in the bottom panel of Fig. 3. While in the case of 78Ni this component is given by $|c_{0}|^{2}\approx 0.48$ indicating a dominant occupation of the energetically lowest configurations in the ground state, this changes drastically towards mid-shell 74Cr with very small $|c_{0}|^{2}\approx 0.03$. Similar observations hold to a lesser extent for the neighboring nuclei 76Fe ($|c_{0}|^{2}\approx 0.28)$ and 72Ti ($|c_{0}|^{2}\approx 0.19$).
@@ 6 Quantum information analysis of ground states
Figure 4: Single-orbital entropies for the ground state of 74Cr in the VS-DMRG, in comparison with increasing CI particle-hole truncations. The corresponding total entropies are given in the legend.
We finally turn to a discussion of orbital entanglement for the strongly correlated 74Cr nucleus using established QI measures. Many-body correlation effects can be explicitly linked to orbital entanglement between single-particle states for a given nuclear wave function [42]. We therefore, define the subsystem’s reduced density matrix $A\subset\mathcal{H}_{1}$ via $\rho_{A}=\text{Tr}_{B}\,\rho_{AB}$, where $A\cup B=\mathcal{H}_{1}$. Setting $A=\{p\}$ we introduce the single-orbital entropy as $s_{i}=-\text{Tr}\,\rho\,\log\,\rho$ that is directly linked to the occupation numbers $n_{p}$ via $s_{p}=-n_{p}\,\log\,n_{p}-\bar{n}_{p}\,\log\,\bar{n}_{p}$ where $\bar{n}_{p}\equiv 1-n_{p}$. Therefore, single-orbital entropies can be straightforwardly evaluated in the DMRG and CI frameworks. Our previous work showed that the orbital entropies can be linked to the closed-shell character of a nucleus [43].
@@ 7 Conclusion and outlook
Finally our methodology was applied to $N=50$ isotones to obtain an ab initio description of transitional nuclei at the $N=50$ shell closure. While our predictions qualitatively agree with previous state-of-the-art shell-model calculations, we fail to reproduce the deformed excited-state band in 78Ni, hinting at increased many-body uncertainties or interaction deficiencies.
While the computational demands of the VS-DMRG implementation are tractable in this work, we envision more advanced schemes to cope with computational challenges emerging in larger valence spaces. This can be either achieved by a symmetry-adapted formulation of the DMRG explicitly accounting for non-Abelian SU(2) symmetries (see Ref. [71]), or by a more careful subspace selection using restricted-active-space formulations, where the coupling to weakly entangled high-energy orbits is approximated by limiting oneself to low-rank CI excitations [72, 61]. This work paves the way for accessing configuration spaces well above current shell-model capacities. These arise for heavy nuclei and nuclei far away from shell closures, and will be particularly important for deriving ab initio constraints on heavy neutron-rich nuclei.
@@ Acknowledgements
The work of A.T., T.M., and A.S. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842). M.A.W. was supported by the Bolyai Research Scholarship of the Hungarian Academy of Sciences. Ö.L. and G.Z. were supported by the Quantum Information National Laboratory of Hungary and by the Hungarian National Research Development and Innovation Office (NKFIH) through Grants Nos. K134983, SNN139581, and TKP2021-NVA-04. Ö.L. received further support from the Hans Fischer Senior Fellowship programme funded by the Technical University of Munich – Institute for Advanced Study and from the Center for Scalable and Predictive methods for Excitation and Correlated phenomena (SPEC), funded as part of the Computational Chemical Sciences Program FWP 70942 by the U.S. Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences, and Biosciences at Pacific Northwest National Laboratory. Computations were in part performed with an allocation of computing resources at the Jülich Supercomputing Center. We have used the NuHamil [73], imsrg++ [74], and KSHELL [28] codes for the 1.8/2.0 (EM) NN+3N matrix elements generation, VS-IMSRG decoupling, and CI diagonalization, respectively.